OpenStudy (anonymous):

Maya plants 8 flower bulbs in the fall: 2 orange,3 pink,and 3 red. Find the probability that the first flower to bloom is orange

OpenStudy (anonymous):

2/8=1/4=25%

OpenStudy (anonymous):

is it 1/4?

OpenStudy (anonymous):

probability is simply the number of outcomes that fit what you want divided by the total possible outcomes.